OpenStudy (anonymous):

Find a • b. a = 4i + 8j, b = -2i + 3j

OpenStudy (amistre64):

thought that was a minus sign

OpenStudy (amistre64):

looks like you want to do a dot product?

OpenStudy (anonymous):

<2, 11> 32 <-8, 24> 16 These are my choices

OpenStudy (anonymous):

Im taking a quiz and I cant figure this one out.

OpenStudy (danjs):

\[A * B = AxBx + AyBy \]

OpenStudy (amistre64):

then you need to refresh yourself on the dot product operation.

OpenStudy (danjs):

The dot product results in a scalar quantity, not a vector

OpenStudy (anonymous):

wouldnt it be C

OpenStudy (amistre64):

the vectors are written in a different notation is all: if u = (a,b) then we can rewrite it as: u = ai + bj thats all

OpenStudy (danjs):

No C is a vector, the dot product results in a number only

OpenStudy (danjs):

Ax*Bx + Ay*By

OpenStudy (anonymous):

so 32

OpenStudy (danjs):

4(-2) + 8(3) =...

OpenStudy (anonymous):

forgot about the negative thank you @DanJS

OpenStudy (danjs):

no prob, remember for vector multiplication, Dot Product = Scalar Cross Product = vector orthogonal to both A and B and has a magnitude of the area spanned by A and B

OpenStudy (anonymous):

thanks again